# Task-Augmented Cross-View Imputation Network for Partial Multi-View Incomplete Multi-Label Classification

### v1

@@ IV-B Experimental Result and Analysis
Corel5k 1-RL 0.891$\pm$0.002 0.926$\pm$0.001 0.884$\pm$0.002 0.890$\pm$0.002 0.929$\pm$0.001 0.930$\pm$0.002 0.949$\pm$0.003 0.953$\pm$0.002
AUC 0.895$\pm$0.002 0.929$\pm$0.001 0.887$\pm$0.002 0.893$\pm$0.002 0.931$\pm$0.000 0.930$\pm$0.003 0.951$\pm$0.002 0.955$\pm$0.002
AP 0.610$\pm$0.005 0.589$\pm$0.006 0.468$\pm$0.004 0.496$\pm$0.003 0.608$\pm$0.003 0.652$\pm$0.008 0.681$\pm$0.007 0.698$\pm$0.009
1-HL 0.926$\pm$0.000 0.931$\pm$0.001 0.888$\pm$0.001 0.929$\pm$0.001 0.937$\pm$0.001 0.939$\pm$0.001 0.943$\pm$0.006 0.944$\pm$0.001
Pascal07 1-RL 0.866$\pm$0.002 0.873$\pm$0.003 0.763$\pm$0.007 0.795$\pm$0.001 0.859$\pm$0.003 0.890$\pm$0.004 0.904$\pm$0.004 0.910$\pm$0.005
AUC 0.879$\pm$0.002 0.893$\pm$0.002 0.793$\pm$0.005 0.822$\pm$0.001 0.876$\pm$0.002 0.898$\pm$0.004 0.914$\pm$0.004 0.918$\pm$0.004
AP 0.264$\pm$0.002 0.392$\pm$0.002 0.260$\pm$0.004 0.251$\pm$0.002 0.391$\pm$0.001 0.405$\pm$0.004 0.373$\pm$0.005 0.410$\pm$0.005
1-HL 0.983$\pm$0.000 0.983$\pm$0.000 0.972$\pm$0.000 0.983$\pm$0.000 0.984$\pm$0.000 0.983$\pm$0.000 0.983$\pm$0.000 0.984$\pm$0.000
ESPGame 1-RL 0.804$\pm$0.003 0.875$\pm$0.001 0.817$\pm$0.001 0.825$\pm$0.001 0.871$\pm$0.003 0.884$\pm$0.002 0.888$\pm$0.002 0.898$\pm$0.002
AUC 0.810$\pm$0.003 0.878$\pm$0.001 0.822$\pm$0.001 0.830$\pm$0.002 0.874$\pm$0.004 0.880$\pm$0.002 0.891$\pm$0.001 0.901$\pm$0.001
AP 0.330$\pm$0.003 0.424$\pm$0.003 0.250$\pm$0.004 0.267$\pm$0.001 0.418$\pm$0.006 0.443$\pm$0.005 0.451$\pm$0.004 0.470$\pm$0.005
1-HL 0.980$\pm$0.000 0.982$\pm$0.000 0.969$\pm$0.000 0.980$\pm$0.000 0.982$\pm$0.000 0.981$\pm$0.000 0.982$\pm$0.000 0.982$\pm$0.000
IAPRTC12 1-RL 0.871$\pm$0.002 0.905$\pm$0.001 0.842$\pm$0.002 0.854$\pm$0.001 0.912$\pm$0.003 0.923$\pm$0.003 0.930$\pm$0.002 0.934$\pm$0.002
AUC 0.877$\pm$0.002 0.906$\pm$0.001 0.843$\pm$0.002 0.855$\pm$0.001 0.911$\pm$0.003 0.922$\pm$0.003 0.931$\pm$0.002 0.935$\pm$0.002
AP 0.643$\pm$0.003 0.665$\pm$0.003 0.493$\pm$0.022 0.555$\pm$0.002 0.659$\pm$0.004 0.683$\pm$0.004 0.701$\pm$0.003 0.712$\pm$0.004
1-HL 0.874$\pm$0.001 0.896$\pm$0.001 0.840$\pm$0.004 0.882$\pm$0.001 0.902$\pm$0.001 0.909$\pm$0.001 0.910$\pm$0.001 0.911$\pm$0.001
MIRFLICKR 1-RL 0.876$\pm$0.006 0.894$\pm$0.001 0.806$\pm$0.015 0.847$\pm$0.001 0.895$\pm$0.001 0.910$\pm$0.002 0.916$\pm$0.001 0.920$\pm$0.001
AUC 0.869$\pm$0.005 0.880$\pm$0.001 0.789$\pm$0.022 0.839$\pm$0.001 0.876$\pm$0.002 0.894$\pm$0.004 0.898$\pm$0.001 0.901$\pm$0.001
(a) Corel5k
(b) Pascal07
@@ IV-C Hyper-parameter Analysis
IV-D Ablation Experiments
In the ablation study, the performance of each TACVI-Net module is evaluated on three representative datasets. The experiment is conducted under the conditions where 50% of the instances from each view are utilized, with an additional 50% of these instances having missing labels, and 70% of the samples for training. To delve deeper into the effectiveness of individual components, a dual-pronged approach is adopted. Initially, two separable loss functions, $L_{ta}^{l}$ and $L_{re}$, are removed from the foundational model to establish a baseline. Subsequently, each of these losses is reintroduced incrementally to the baseline, allowing for a meticulous examination of their individual contributions to enhancing model performance. In parallel, to comprehensively gauge the efficacy of data imputation strategies, a contrasting experiment is carried out where no data imputation was applied. This direct comparison serve to highlight the profound impact that imputation techniques have on the overall model performance. The relevant experimental results are summarized in Table V. We can observe that each component of TACVI-Net is vital and beneficial, highlighting the importance of every component in boosting the final classification performance. Furthermore, upon the removal of imputation, there is a noticeable decline in the model’s metrics. This suggests that the imputation strategy indeed addresses issues associated with insufficient information.
TABLE V: Ablation experimental results of TACVI on the Corel5k, Pascal07 and ESPGame datasets with 50% missing views, 50% missing labels, and 70% training samples.
@@ IV-D Ablation Experiments
without imputaion 0.392 0.907 0.560 0.855 0.305 0.848
V Conclusion
In this paper, we innovatively introduce a task-augmented cross-view imputation network (TACVI-Net), which departs from previous approaches that rely on assumed missing information to bypass incomplete views. Departing from such conventions, TACVI-Net employs a novel dual-stage strategy which limits the accumulation of redundant information while filling in the missing data so that the model has enough and useful information to learn inter-view consistency and complementarity to improve classification performance. Extensive experiments on five datasets confirm TACVI-Net’s superior performance and highlight its innovative capabilities in addressing the challenges of partial multi-view incomplete multi-label classification task.

### v2

@@ 4.2 Experimental Result and Analysis
Corel5k 1-RL 0.891$\pm$0.002 0.926$\pm$0.001 0.884$\pm$0.002 0.890$\pm$0.002 0.929$\pm$0.001 0.931$\pm$0.001 0.949$\pm$0.003 0.953$\pm$0.002
AUC 0.895$\pm$0.002 0.929$\pm$0.001 0.887$\pm$0.002 0.893$\pm$0.002 0.931$\pm$0.000 0.934$\pm$0.001 0.951$\pm$0.002 0.955$\pm$0.002
AP 0.610$\pm$0.005 0.589$\pm$0.006 0.468$\pm$0.004 0.496$\pm$0.003 0.608$\pm$0.003 0.648$\pm$0.006 0.681$\pm$0.007 0.698$\pm$0.009
1-HL 0.926$\pm$0.000 0.931$\pm$0.001 0.888$\pm$0.001 0.929$\pm$0.001 0.937$\pm$0.001 0.940$\pm$0.001 0.943$\pm$0.006 0.944$\pm$0.001
Pascal07 1-RL 0.866$\pm$0.002 0.873$\pm$0.003 0.763$\pm$0.007 0.795$\pm$0.001 0.859$\pm$0.003 0.886$\pm$0.003 0.904$\pm$0.004 0.910$\pm$0.005
AUC 0.879$\pm$0.002 0.893$\pm$0.002 0.793$\pm$0.005 0.822$\pm$0.001 0.876$\pm$0.002 0.897$\pm$0.003 0.914$\pm$0.004 0.918$\pm$0.004
AP 0.264$\pm$0.002 0.392$\pm$0.002 0.260$\pm$0.004 0.251$\pm$0.002 0.391$\pm$0.001 0.397$\pm$0.002 0.373$\pm$0.005 0.410$\pm$0.005
1-HL 0.983$\pm$0.000 0.983$\pm$0.000 0.972$\pm$0.000 0.983$\pm$0.000 0.984$\pm$0.000 0.984$\pm$0.000 0.983$\pm$0.000 0.984$\pm$0.000
ESPGame 1-RL 0.804$\pm$0.003 0.875$\pm$0.001 0.817$\pm$0.001 0.825$\pm$0.001 0.871$\pm$0.003 0.880$\pm$0.001 0.888$\pm$0.002 0.898$\pm$0.002
AUC 0.810$\pm$0.003 0.878$\pm$0.001 0.822$\pm$0.001 0.830$\pm$0.002 0.874$\pm$0.004 0.883$\pm$0.001 0.891$\pm$0.001 0.901$\pm$0.001
AP 0.330$\pm$0.003 0.424$\pm$0.003 0.250$\pm$0.004 0.267$\pm$0.001 0.418$\pm$0.006 0.443$\pm$0.003 0.451$\pm$0.004 0.470$\pm$0.005
1-HL 0.980$\pm$0.000 0.982$\pm$0.000 0.969$\pm$0.000 0.980$\pm$0.000 0.982$\pm$0.000 0.982$\pm$0.000 0.982$\pm$0.000 0.982$\pm$0.000
IAPRTC12 1-RL 0.871$\pm$0.002 0.905$\pm$0.001 0.842$\pm$0.002 0.854$\pm$0.001 0.912$\pm$0.003 0.918$\pm$0.002 0.930$\pm$0.002 0.934$\pm$0.002
AUC 0.877$\pm$0.002 0.906$\pm$0.001 0.843$\pm$0.002 0.855$\pm$0.001 0.911$\pm$0.003 0.918$\pm$0.002 0.931$\pm$0.002 0.935$\pm$0.002
AP 0.643$\pm$0.003 0.665$\pm$0.003 0.493$\pm$0.022 0.555$\pm$0.002 0.659$\pm$0.004 0.681$\pm$0.004 0.701$\pm$0.003 0.712$\pm$0.004
1-HL 0.874$\pm$0.001 0.896$\pm$0.001 0.840$\pm$0.004 0.882$\pm$0.001 0.902$\pm$0.001 0.906$\pm$0.001 0.910$\pm$0.001 0.911$\pm$0.001
MIRFLICKR 1-RL 0.876$\pm$0.006 0.894$\pm$0.001 0.806$\pm$0.015 0.847$\pm$0.001 0.895$\pm$0.001 0.907$\pm$0.002 0.916$\pm$0.001 0.920$\pm$0.001
AUC 0.869$\pm$0.005 0.880$\pm$0.001 0.789$\pm$0.022 0.839$\pm$0.001 0.876$\pm$0.002 0.888$\pm$0.002 0.898$\pm$0.001 0.901$\pm$0.001
4.3 Hyper-parameter Analysis
@@
4.4 Ablation Experiments
In the ablation study, the performance of each TACVI-Net module is evaluated on three representative datasets. The experiment is conducted under the conditions where 50% of the instances from each view are utilized, with an additional 50% of these instances having missing labels, and 70% of the samples for training. To delve deeper into the effectiveness of individual components, a dual-pronged approach is adopted. Initially, two separable loss functions, $L_{ta}^{l}$ and $L_{re}$, are removed from the foundational model to establish a baseline. Subsequently, each of these losses is reintroduced incrementally to the baseline, allowing for a meticulous examination of their individual contributions to enhancing model performance. In parallel, to comprehensively gauge the efficacy of data imputation strategies, a contrasting experiment is carried out where no data imputation was applied. This direct comparison serve to highlight the profound impact that imputation techniques have on the overall model performance. The relevant experimental results are summarized in Table 5. We can observe that each component of TACVI-Net is vital and beneficial, highlighting the importance of every component in boosting the final classification performance. Furthermore, upon the removal of imputation, there is a noticeable decline in the model’s metrics. This suggests that the imputation strategy indeed addresses issues associated with insufficient information.
Table 5: Ablation experimental results of TACVI-Net on the Corel5k, Pascal07 and ESPGame datasets with 50% missing views, 50% missing labels, and 70% training samples.
@@ 4.4 Ablation Experiments
without imputation 0.392 0.907 0.560 0.855 0.305 0.848
5 Conclusion
In this paper, we innovatively introduce a task-augmented cross-view imputation network (TACVI-Net), which departs from previous approaches that rely on assumed missing information to bypass incomplete views. Departing from such conventions, TACVI-Net employs a novel dual-stage strategy which limits the accumulation of redundant information while filling in the missing data so that the model has enough and useful information to learn inter-view consistency and complementarity to improve classification performance. Extensive experiments on five datasets confirm TACVI-Net’s superior performance and highlight its innovative capabilities in addressing the challenges of partial multi-view incomplete multi-label classification task. However, our method does not fully consider the differences in quality among different views. Therefore, future research will more deeply explore how to effectively assess and utilize the quality of each view during the fusion process, to further enhance the robustness and overall performance of multi-view learning.
6 Acknowledgements
This work is supported by Guizhou University Basic Research Program ([2024] 04) and Guizhou Provincial Basic Research Program (Natural Science) (No. QianKeHe Basic - [2024] Youth 093).